OpenStudy (anonymous):

sin theta = 1/9 and cot theta= 36/13 what is cos theta

OpenStudy (lgbasallote):

sin is opposite/hypotenuse cot is adjacent/opposite cos is adjacent/ hypotenuse

OpenStudy (lgbasallote):

ring a bell?

OpenStudy (anonymous):

Sin=opp/hyp. Cos=adj/hyp

OpenStudy (anonymous):

Draw a triangle with your values. Theta is your angles and you have opposite and adjacent values.

OpenStudy (anonymous):

cos theta = 36/9?

OpenStudy (lgbasallote):

yes!!

OpenStudy (anonymous):

do you have to simplify?

OpenStudy (lgbasallote):

cot = 36/13? is it really over 13?

OpenStudy (anonymous):

Draw a triangle with your values. Theta is your angles and you have opposite and adjacent values.

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

then you have to make 1/9 have a numerator of 13 sin = 13/117 therefore hypotenuse is 117

OpenStudy (lgbasallote):

therefore it isnt 36/9 sorry...

OpenStudy (anonymous):

36/117 then, so why isn't it 36/9?

OpenStudy (lgbasallote):

because it says in cot that 36/13 therefore 13 should be the opposite but in sin it says that sin = opposite/hypotenuse = 1/9 therefore we turn the opposite into 13

OpenStudy (anonymous):

ooo ok so is cos 36/117 then